OpenStudy (anonymous):

An Air Force C-130 Hercules cargo plane left Sheppard Air Force Base in Wichita Falls, TX at a height of 3,958 feet above sea level. If the plane climbs at a rate of 389 feet per minute, which function could be used to determine t, the time in minutes it will take the plane to reach its cruising elevation of 12,127 feet above sea level?

OpenStudy (anonymous):

12,127 = 389 + 3,958t 12,127 = 389(t + 3,958) 12,127 = 3,958 + 389t 12,127 = (3,958 + 389)t possible answers